Skip to main contentSkip to solution

A doctor is to visit a patient. From the past experiences, it is known that the probabilities of his coming by cab, metro, bike or by other means of transport are respectively 0.3, 0.2, 0.1 and 0.4 respectively. The probabilities that he will be late are 0.25, 0.3, 0.35 and 0.1 if he comes by cab, metro, bike and by other means of transport respectively. using this information about the doctor, answer the following questions :

When the doctor arrives late, what is the probability that he comes by metro?

Solution

✅ Correct Option: 2

Total probability of arriving late:

P(L)=0.3(0.25)+0.2(0.3)+0.1(0.35)+0.4(0.1)=0.075+0.06+0.035+0.04=0.21P(L) = 0.3(0.25) + 0.2(0.3) + 0.1(0.35) + 0.4(0.1) = 0.075 + 0.06 + 0.035 + 0.04 = 0.21.

By Bayes' theorem:

P(metro∣L)=0.2×0.30.21=0.060.21=621=27P(\text{metro}|L) = \frac{0.2 \times 0.3}{0.21} = \frac{0.06}{0.21} = \frac{6}{21} = \frac{2}{7}.

Keyboard Shortcuts

  • Left arrow: Previous question
  • Right arrow: Next question
  • S key: Jump to solution
  • Q key: Jump to question